OpenStudy (anonymous):

f(x)=(x+6)^2+2 finding the roots????

OpenStudy (anonymous):

expand. set the function equal to zero and quadraticize the function.

OpenStudy (anonymous):

How do you do that exactly? What does expand it mean? and to quadraticize?

OpenStudy (anonymous):

f(x) = (x+6)^2+2 f(x) = x^2+12x+36+2 f(x) = x^2+12x+38 // either graph this function, or set f(x) = 0 and use the quadratic formula. //quadratic formula method x^2+12x+38 = 0 to find the roots use the quadratic. if you don't know what that is, look on Khan Academy.